OpenStudy (anonymous):

Matthews 4th grade science experiment counting the number of popcorn kernels that popped after 3 minutes in a microwave.if 126 kernels popped and 38 kernels did not pop what was the percent of kernels popped to the nearest whole percent??

OpenStudy (beccaboo333):

I'm not entirely positive. I think around 30% to give you an idea

OpenStudy (beccaboo333):

@ganeshie8

OpenStudy (anonymous):

who that

OpenStudy (beccaboo333):

A math genius O_O

OpenStudy (anonymous):

ok and thx

ganeshie8 (ganeshie8):

lol okay, total number of kernels = 126 + 38 = 164 number of kernels popped = 126 you need to find out what percent of 164 is 126

OpenStudy (anonymous):

ok thx

ganeshie8 (ganeshie8):

do u knw how to find the percentage ha ? :)

OpenStudy (anonymous):

noo

ganeshie8 (ganeshie8):

here is how to find it :- \(\large \frac{126}{164} \times 100 \)

OpenStudy (anonymous):

76.82926829

OpenStudy (anonymous):

is tht correct

OpenStudy (beccaboo333):

@ganeshie8 COME BACK

OpenStudy (beccaboo333):

My guess it to take 100 and subtract 76.82926829 from it to get your full answer of how many didn't pop

ganeshie8 (ganeshie8):

76.8 % of kernels popped (100-76.8) % of kernels didnt pop

ganeshie8 (ganeshie8):

but, the question is asking for % of kernels popped so answer wud be 76.8%

OpenStudy (anonymous):

ok thx

OpenStudy (beccaboo333):

oh. whoops lol.. see why @ganeshie8 is the genius xD

OpenStudy (anonymous):

yup